OpenStudy (shamil98):

Could I get an explanation of the power rule on derivatives? .. For example; 4x^3 - 5x^1/3 +3x^-2

OpenStudy (abb0t):

Suppose you have: \(\sf \color{red}{ax^{n}}\), if youre trying to find the derivative: \(\sf \color{}{\frac{d}{dx}(ax^n) = nax^{n-1}}\)

OpenStudy (shamil98):

Oh, so I know 4x^3 becomes 12x^2

OpenStudy (abb0t):

So for the first one, you have 4x\(^3\), it would be: \(\sf \color{blue}{3 \times 4x^{3-1}}\)

OpenStudy (abb0t):

Yeah, 12x\(^2\)

OpenStudy (shamil98):

What would -5x^1/3 and 3x^-2 become?..

OpenStudy (abb0t):

\(\sf \color{red}{\frac{1}{3}(5)x^{\frac{1}{3}-1}}\)

OpenStudy (shamil98):

ah, so 5/3 x^ -2/3 ?..

OpenStudy (shamil98):

But you can't have a negative exponent..

OpenStudy (abb0t):

Why not?

OpenStudy (shamil98):

Aren't you suppose to put something like x^-2 into 1/x^2

OpenStudy (shamil98):

or do i just leave it as that?.

OpenStudy (abb0t):

\(\sf \color{}{x^{-\frac{2}{3}}=\frac{1}{x^{\frac{2}{3}}}}\) or \(\frac{ 1 }{ \sqrt[3]{x^2} }\)

OpenStudy (shamil98):

Alright, so would 3x^-2 become -6/x ?

OpenStudy (abb0t):

No. -2-1 = -1?

OpenStudy (shamil98):

so just -6x^-1 then..

OpenStudy (abb0t):

\(\huge -2-1 = -1?\) really??

OpenStudy (shamil98):

oh oops, -3 LOL

OpenStudy (shamil98):

still morning forgive me .. >.>

OpenStudy (abb0t):

You're such a noob. I'm done with you. Good bye.

OpenStudy (shamil98):

LOL

OpenStudy (shamil98):

alright thanks :P